a car changes its speed from 60 m/s to 40 m/s it does this at a rat of four seconds what was the acceleration of the car
1
Expert's answer
2012-05-24T08:46:17-0400
V0 = 60 m/s
V1 = 40 m/s
dt = 4 s
a = dV / dt = (V1 - V0) / dt = (40 -
60) / 4 = -5 m / s^2